OpenStudy (help_people):

Given the following functions f(x) and g(x), solve f[g(7)] and select the correct answer below: f(x) = 4x + 21 g(x) = 2x + 2 16 49 85 784

Vocaloid (vocaloid):

g(x) = 2x + 2 g(7) = ?

OpenStudy (help_people):

g(7)=2(7)+2 g(7)=14+2 g(7)=16

OpenStudy (help_people):

@Vocaloid

Vocaloid (vocaloid):

right, now on to the next step f(x) = 4x + 21 f(g(7)) = ?

OpenStudy (help_people):

not sure :(

Vocaloid (vocaloid):

g(7) = 16 f(g(7)) = f(16) = ?

OpenStudy (help_people):

wel dint we have to find f(X) ?

OpenStudy (help_people):

i don't get this part

Vocaloid (vocaloid):

please look at the problem what is f(x)?

OpenStudy (help_people):

f(x)=4x+21

Vocaloid (vocaloid):

right. f(x)=4x+21 f(g(7)) = f(16) = ?

OpenStudy (help_people):

f(x)=4(16)+21 ?

Vocaloid (vocaloid):

yes

OpenStudy (help_people):

ok

OpenStudy (help_people):

64 +21=85

Vocaloid (vocaloid):

right, that's our answer

OpenStudy (help_people):

ty!!! May you help w. ones similar to those ?

OpenStudy (help_people):

i will post on another tab :)

Vocaloid (vocaloid):

sure

OpenStudy (help_people):

:)